# references.bib

### v3

@@ Abstract
The era of vision-language models (VLMs) trained on large web-scale datasets challenges conventional formulations of “open-world" perception. In this work, we revisit the task of few-shot object detection (FSOD) in the context of recent foundational VLMs. First, we point out that zero-shot VLMs such as GroundingDINO significantly outperform state-of-the-art few-shot detectors (48 vs. 33 AP) on COCO. Despite their strong zero-shot performance, such foundational models may still be sub-optimal. For example, trucks on the web may be defined differently from trucks for a target application such as autonomous vehicle perception. We argue that the task of few-shot recognition can be reformulated as aligning foundation models to target concepts using a few examples. Interestingly, such examples can be multi-modal, using both text and visual cues, mimicking instructions that are often given to human annotators when defining a target concept of interest. Concretely, we propose Foundational FSOD, a new benchmark protocol that evaluates detectors pre-trained on any external datasets and fine-tuned on multi-modal (text and visual) $K$-shot examples per target class. We repurpose nuImages for Foundational FSOD, benchmark several popular open-source VLMs, and provide an empirical analysis of state-of-the-art methods. Lastly, we discuss our recent CVPR 2024 Foundational FSOD competition and share insights from the community. Notably, the winning team significantly outperforms our baseline by 23.9 mAP!
@@ 1 Introduction
Vision-language models (VLMs) trained on (often proprietary) web-scale datasets have disrupted traditional notions of the "open-world," particularly for few-shot recognition. In this paper, we revisit few-shot object detection (FSOD) in the context of these foundational models, propose a new benchmark protocol that allow foundational models to “enter the conversation", and present several simple baselines.
First, we highlight that zero-shot VLMs like GroundingDINO demonstrate a remarkable improvement over state-of-the-art few-shot detectors (48.3 vs. 33.1 AP) on COCO, as shown in Table 1. In hindsight, this is not surprising, as the former is pre-trained on far more data (that may include visual examples of the target concept), while the later is pre-trained on data that is explicitly curated to avoid target concepts of interest. From this perspective, VLMs violate the current training protocol of few-shot benchmarks, suggesting that such protocols need to be rethought in the foundational era.
@@
Figure 1: Poor Alignment Between Vision Language Models (VLMs) and Target Concepts. Although VLMs show impressive zero-shot performance, they struggle when the target class is different from concepts encountered in web-scale training. On the left, we see that the nuImages dataset [caesar2020nuscenes] defines the cab of the truck as a separate concept from its trailer (shown in red). In contrast, the VLM predicts the entire vehicle as a truck (shown in green). Similarly, nuImages annotations dictate that a person riding a bicycle must also be labeled as part of bicycle (shown in red) unlike the VLM prediction (in green). On the right, we present the actual class definitions given to the nuImages annotators, provided as both textual descriptions and visual examples. Just as human annotators learn concepts from few-shot multi-modal examples, we argue that VLMs should be aligned to $K$ vision-language examples.
@@
Conventional FSOD. Existing FSOD benchmarks partition object detection datasets like PASCAL VOC [Everingham10] and COCO [Lin2014MicrosoftCC] into base and novel classes. Detectors pre-train on base and then learn novel classes given $K$ examples (or $K$-shots). Current protocols enforce base and novel to be disjoint to prevent concept leakage, allowing one to evaluate generalization to the “open-world". However, as most detectors are pre-trained on ImageNet, we point out that concept leakage already occurs in current FSOD protocols. For example, cat and person are deemed novel for COCO-FSOD but are present in ImageNet data used to pre-train detectors [wang2020frustratingly]. Moreoever, car is deemed novel, but similar concepts like sports car and race car are present in ImageNet, illustrating the difficulty of even defining leakage.
@@
Foundational FSOD. We believe that concept leakage should be embraced. Our Foundational FSOD protocol replaces the base pre-training stage with web-scale pre-training, where such data may be proprietary and not fully disclosed [radford2021learning]. We argue that pre-training on large-scale data will be the key enabler for generalization to the open-world. Note that this hypothesis is difficult to even evaluate under current few-shot protocols, motivating our setup. Moreover, another key property is that $K$-shots may include multi-modal examples spanning both images and text, motivating a multi-modal adaptation stage that aligns the VLM to the target concepts (cf. Fig. 2). We repurpose nuImages for our Foundational FSOD benchmark, a challenging dataset due to open-world categories such as debris and pushable-pullable, which also provides multi-modal annotation instructions.
Contributions. We present three major contributions. First, we modernize FSOD benchmarks by embracing vision-language foundation models that are pretrained on internet-scale data. We highlight the practical challenge of using multi-modal few-shot examples to define the target semantic concept (as shown in Fig. 1). Next, we adapt nuImages for Foundational FSOD, evaluate various popular open-source VLMs, and present an empirical analysis of leading methods. Lastly, we highlight the results from our recent CVPR 2024 challenge hosted in conjunction with the Workshop on Visual Perception via Learning in An Open World.
@@
Figure 2: Foundational Few-shot Object Detection (FSOD). Conventional FSOD protocols (left) allow for pre-training on base classes (with many examples per class) and then fine-tuning on $K$-shots of novel classes, where novel and base are designed to be disjoint. However, we point out that pre-training datasets such as ImageNet often contain classes similar to the novel classes, highlighting the issue of concept leakage. As preventing concept leakage is difficult (if not impossible) and appears to be artificial in the foundational era, we propose the setup of foundational FSOD (right). Our setup allows for pre-training on massively-large (and potentially proprietary) datasets, typical for foundational VLM models. Since these models can process both text and images, one can utilize such multi-modal $K$-shot examples to align VLMs with the target concepts of interest.
@@ 2 Related Works
Few-Shot Object Detection aims to detect new categories with limited training data [kohler2021few]. Recent work explores two primary approaches: meta-learning and transfer learning. Meta-learning-based methods focus on acquiring generalizable features from a set of base classes, which can then be applied to identify novel classes. For example, [kang2019few] proposes a technique that re-weights features from base classes to predict novel classes. [xiao2022few] proposes a framework addressing both few-shot object detection and few-shot viewpoint estimation. [fan2020few] introduces a general FSOD network that learns a matching metric between image pairs, while [wu2021universal] enhances object features using a universal prototype. More recently, [xu2023generating] proposes a generative approach that is robust to noisy object proposals for novel classes. In contrast, transfer learning involves partially freezing network weights pretrained on a base dataset to improve a model’s ability to generalize to novel classes with limited data. Transfer learning approaches often follow a two-stage fine-tuning strategy: first train on base classes and then fine-tune the box classifier and regressor with $K$-shots from novel classes. This strategy has historically outperformed meta-learning approaches [wang2020frustratingly]. Recent work has primarily focused on improving classification performance. [sun2021fsce] utilizes a contrastive proposal encoding loss to encourage instance-level intra-class compactness and inter-class variance. Similarly, [li2021beyond] applies a class margin loss to balance inter and intra-class margins. Our approach leverages transfer-learning by fine-tuning vision-language models (VLMs) pre-trained on large-scale datasets.
@@
Vision Language Models are trained on a large-scale collection of weakly-supervised image-text pairs collected from the web. These models embed images and text into a shared space, enabling open-vocabulary detection. Early works adapt VLMs for object detection by either distilling the model’s predictions for specific image regions [gu2021vild, gu2021open] or directly incorporating detection components into frozen [kuo2022fvlm] or fine-tuned [minderer2022owlvit, minderer2023owlvit2, du2022learning] encoders. In contrast, RegionCLIP [zhong2022regionclip] employs a multi-stage training approach, which involves generating pseudo-labels from captioning data, conducting region-text contrastive pre-training, and fine-tuning on detection data. GLIP [li2021grounded] uses a single text query for the entire image and frames detection as a phrase grounding problem. Detic [zhou2022detecting] addresses long-tail detection performance by leveraging image-level supervision. In the context of open-vocabulary detection, there may be some overlap between categories seen during training and testing. We use the term “zero-shot inference” to signify that a model has never been trained on the target dataset.
@@
Fine-Tuning Foundation Models is of significant interest across many application areas [hu2021lora, zhang2023adding, gao2024clip]. Standard fine-tuning procedures employ both linear probing [chen2020simple, he2022masked, he2020momentum] and full-finetuning [wang2017growing, wu2023revisiting, kirkpatrick2017overcoming] to adapt models to downstream tasks. However, such methods may be suboptimal as they can be computationally expensive. Instead, recent works like CLIP-Adapter [gao2024clip] and Tip-Adapter [zhang2021tip] fine-tune CLIP using parameter-efficient methods [houlsby2019parameter, zhang2020side, jia2022visual] which optimize lightweight MLPs while keeping the encoder frozen. Similarly, inspired by the success of prefix-tuning in language models [deng2022rlprompt, jiang2020can, haviv2021bertese, gao2020making], prompt adaptation [lu2022prompt, zhu2023prompt, xing2023dual, zhou2022conditional] replaces hand-crafted prompts like "a photo of a {cls}" with learned word embeddings. CoOp [zhou2022learning] and other prompting methods [lu2022prompt, zhu2023prompt, zhou2022conditional] finetune CLIP via prefix-tuning. Different from most prior work, we investigate fine-tuning strategies for VLM-based detectors using few-shot multi-modal examples.
@@ 3.1 Foundational FSOD Benchmark
Existing FSOD benchmarks repurpose well-established datasets like PASCAL VOC [Everingham10] and COCO [Lin2014MicrosoftCC] by partitioning them into base and novel classes for pre-training and fine-tuning, respectively. For COCO, the 60 categories disjoint with PASCAL VOC are used as base classes and the remaining 20 are used as novel classes [wang2020frustratingly]. However, this setup is artificial and does not reflect how FSOD is deployed in practice. First, the FSOD benchmarks construct novel classes by including common concepts such as car and person, and require FSOD methods to detect these common classes by assuming they have only few examples. Importantly, VLMs like GroundingDINO [liu2023grounding] can already detect common categories with high accuracy on COCO without fine-tuning (cf. Table 1). Therefore, we focus on benchmarking Foundational FSOD on more realistic and challenging datasets like nuImages [caesar2020nuscenes]. Second, existing FSOD benchmarks require that datasets are partitioned into base and novel classes, which is infeasible for large-scale (often private) foundational datasets. For example, although CLIP’s [radford2021learning] model weights are publicly available, its pre-training dataset is not. Instead, FSOD methods should only fine-tune VLMs on $K$-shot annotations for $C$ target classes (or novel, as termed in the conventional FSOD benchmark), and also evaluate performance on these $C$ classes.
@@ 3.2 Few-Shot Multi-Modal Concept Alignment
Although VLMs achieve strong zero-shot performance on common classes, they struggle when the target class is different from concepts encountered on the web (cf. Fig. 1). For example, nuImages [caesar2020nuscenes] defines the cab of a truck as a separate concept from its trailer. However, Detic detects the entire vehicle as truck. This fine-grained distinction is provided to human annotators with visual examples and textual descriptions. We explore seven methods for alignment (either explicitly by updating model weights via fine-tuning or in context via prompting) VLMs below.
Prompt Engineering uses expressive descriptions in the text prompt, adding attributes, synonyms or language context, to manually improve the alignment of foundation model outputs to target concepts of interest. In our case, we prompt VLMs with synonyms of the nuImages classes to improve detection accuracy. For example, we augment the language query for pushable-pullable with synonyms like cart and wheel barrow.
@@
Multi-Modal Prompting combines language and visual prompting to leverage multi-modal features. Intuitively, multi-modal cues can yield better alignment than uni-modal cues alone; in the above case, ambiguous concepts such as debris can be clarified with both textual descriptions (e.g trash can and tree branch) and visual examples (similar to the multi-modal annotator instructions in Fig. 1!). Here, visual and language prompts are tokenized and separately fed as inputs to a frozen VLM. Specifically, MQDet[xu2024multi] introduces a lightweight module: Gated Class Scalable Perceiver, that fuses visual cues and text descriptions in the text encoder via class-wise cross attention layers.
@@ 4 Experiments
We conduct extensive experiments to validate that zero-shot inference from VLMs significantly improves over state-of-the-art FSOD approaches, suggesting that existing benchmarks should be re-framed to allow foundation models to “enter the conversation”. Moreover, we demonstrate that leveraging language cues, especially those available for free (e.g. class names), are crucial to improving performance on data-constrained tasks like FSOD.
Datasets and Metrics. We repurpose nuImages [caesar2020nuscenes] to support the study of Foundational FSOD. This dataset annotates 18 classes, which are divided into groups with many, medium, and few examples [peri2023towards, ma2023longtailed]. We report AP for each cohort. Although this dataset is not traditionally used for FSOD, nuImages’ open-world categories like debris and pushable-pullable make it particularly challenging (even for VLMs), and is a realistic benchmark for Foundational FSOD. We follow the $K$-shot dataset creation process established by [wang2020frustratingly], described below. To construct a $K$-shot dataset, we select a target class $c$ and an image at random. If the total annotations for class $c$ in the image are less than or equal to $K$, we add the image to our dataset. We repeat this process for all classes until we have exactly $K$ annotations per class. Since the specific $K$ examples can have a significant impact on the overall performance, we run each experiment over 3 random data splits and report the average.

### v4

@@ Abstract
The era of vision-language models (VLMs) trained on web-scale datasets challenges conventional formulations of “open-world" perception. In this work, we revisit the task of few-shot object detection (FSOD) in the context of recent foundational VLMs. First, we point out that zero-shot predictions from VLMs such as GroundingDINO significantly outperform state-of-the-art few-shot detectors (48 vs. 33 AP) on COCO. Despite their strong zero-shot performance, such foundation models may still be sub-optimal. For example, trucks on the web may be defined differently from trucks for a target application such as autonomous vehicle perception. We argue that the task of few-shot recognition can be reformulated as aligning foundation models to target concepts using a few examples. Interestingly, such examples can be multi-modal, using both text and visual cues, mimicking instructions that are often given to human annotators when defining a target concept of interest. Concretely, we propose Foundational FSOD, a new benchmark protocol that evaluates detectors pre-trained on any external data and fine-tuned on multi-modal (text and visual) K-shot examples per target class. We repurpose nuImages for Foundational FSOD, benchmark several popular open-source VLMs, and provide an empirical analysis of state-of-the-art methods. Lastly, we discuss our recent CVPR 2024 Foundational FSOD competition and share insights from the community. Notably, the winning team significantly outperforms our baseline by 23.3 mAP! Our code and dataset splits are available on GitHub and HuggingFace.
@@ 1 Introduction
Vision-language models (VLMs) trained on (often proprietary) web-scale datasets have disrupted traditional notions of the “open-world”, particularly for few-shot recognition. In this paper, we revisit few-shot object detection (FSOD) in the context of these foundation models, propose a new benchmark protocol that allows foundation models to “enter the conversation”, and present several simple baselines.
First, we highlight that zero-shot predictions from VLMs like GroundingDINO [liu2023grounding] demonstrate a remarkable improvement over state-of-the-art few-shot detectors (48.3 vs. 33.1 AP) on COCO [Lin2014MicrosoftCC], as shown in Table 1. In hindsight, this is not surprising, as the former is pre-trained on far more data (that may include visual examples of the target concept), while the later is pre-trained on data that is explicitly curated to avoid target concepts of interest. From this perspective, VLMs violate the current training protocol of few-shot benchmarks, suggesting that such protocols need to be rethought in the foundational era.
@@
Figure 1: Poor Alignment Between Vision Language Models (VLMs) and Target Concepts. Although VLMs show impressive zero-shot performance, they struggle when the target class is different from concepts encountered in web-scale training. On the left, we see that the nuImages dataset [caesar2020nuscenes] defines the cab of the truck as a separate concept from its trailer (shown in red). In contrast, the VLM predicts the entire vehicle as a truck (shown in green). Similarly, nuImages annotations dictate that a person riding a bicycle must also be labeled as part of bicycle (shown in red) unlike the VLM prediction (in green). On the right, we present the actual class definitions given to the nuImages annotators, provided as both textual descriptions and visual examples. Just as human annotators learn concepts from few-shot multi-modal examples, we argue that VLMs should be aligned with $K$ vision-language examples.
@@
Conventional FSOD. Existing FSOD benchmarks partition object detection datasets like PASCAL VOC [Everingham10] and COCO [Lin2014MicrosoftCC] into base and novel classes. Detectors pre-train on base classes and then learn to identify novel classes given $K$ examples (or $K$-shots). Current protocols enforce base and novel to be disjoint to prevent concept leakage, allowing one to evaluate generalization to the “open-world". However, as most detectors are pre-trained on ImageNet [imagenet], we point out that concept leakage already occurs in the current FSOD protocol. For example, cat and person are deemed novel for COCO-FSOD but are present in ImageNet data used to pre-train detectors [wang2020frustratingly]. Moreoever, car is deemed novel, but similar concepts like sports car and race car are present in ImageNet, illustrating the difficulty of even defining leakage.
@@
Foundational FSOD. We believe that concept leakage should be embraced. Our Foundational FSOD protocol replaces the base pre-training stage with web-scale pre-training, where such data may be proprietary and not fully disclosed [radford2021learning]. We argue that pre-training on large-scale data will be the key enabler for generalization to the open world. Note that this hypothesis is difficult to even evaluate under the conventional few-shot protocol, motivating our setup. Moreover, another key property is that $K$-shot instances may include multi-modal examples spanning both images and text, motivating a multi-modal adaptation stage that aligns the VLM to target concepts (cf. Fig. 2). We repurpose nuImages [caesar2020nuscenes], a challenging dataset due to open-world categories such as debris and pushable-pullable, for our Foundational FSOD benchmark.
We present three major contributions.
  • We modernize FSOD benchmarks by embracing foundational VLMs that are pretrained on internet-scale data. We highlight the practical challenge of using multi-modal few-shot examples to define target semantic concepts (as shown in Fig. 1).
  • We adapt nuImages for Foundational FSOD, evaluate popular open-source VLMs, and present an empirical analysis of leading methods.
  • We highlight the results from our recent CVPR 2024 challenge hosted in conjunction with the Workshop on Visual Perception via Learning in An Open World.
@@
Figure 2: Foundational Few-Shot Object Detection (FSOD). Conventional FSOD protocols (left) allow for pre-training on base classes (with many examples per class) and then fine-tuning on $K$-shots of novel classes, where novel and base are designed to be disjoint. However, we point out that pre-training datasets such as ImageNet often contain classes similar to novel classes, highlighting the issue of concept leakage. As preventing concept leakage is difficult (if not impossible) and appears to be artificial in the foundational era, we propose Foundational FSOD (right). Our setup allows for pre-training on massive (and potentially proprietary) datasets, typical for foundational vision-language models. Since these models can process both text and images, one can utilize such multi-modal $K$-shot examples to align VLMs with the target concepts of interest.
@@ 2 Related Works
Few-Shot Object Detection aims to detect new categories with limited training data [kohler2021few]. Recent work explores two primary approaches: meta-learning and transfer learning. Meta-learning-based methods focus on acquiring generalizable features from a set of base classes, which can then be applied to identify novel classes. For example, Kang et al. [kang2019few] proposes a technique that re-weights features from base classes to predict novel classes. Xiao et al. [xiao2022few] propose a framework addressing both few-shot object detection and few-shot viewpoint estimation. Fan et al. [fan2020few] introduces a general FSOD network that learns a matching metric between image pairs, while Wu et al. [wu2021universal] enhances object features using a universal prototype. More recently, Xu et al. [xu2023generating] propose a generative approach that is robust to noisy object proposals for novel classes. In contrast, transfer learning involves partially freezing network weights pretrained on a base dataset to improve a model’s ability to generalize to novel classes with limited data. Transfer learning approaches often follow a two-stage fine-tuning strategy: first train on base classes and then fine-tune the box classifier and regressor with $K$-shots from novel classes. This strategy has historically outperformed meta-learning approaches [wang2020frustratingly]. Recent work has primarily focused on improving classification performance. Sun et al. [sun2021fsce] utilizes a contrastive proposal encoding loss to encourage instance-level intra-class compactness and inter-class variance. Similarly, Li et al. [li2021beyond] applies a class margin loss to balance inter and intra-class margins.
@@
Vision Language Models are trained on a large-scale collection of weakly-supervised image-text pairs collected from the web. These models embed images and text into a shared space, enabling open-vocabulary detection. Early works adapt VLMs for object detection by either distilling the model’s predictions for specific image regions [gu2021vild, gu2021open] or directly incorporating detection components into frozen [kuo2022fvlm] or fine-tuned [minderer2022owlvit, minderer2023owlvit2, du2022learning] encoders. In contrast, RegionCLIP [zhong2022regionclip] employs a multi-stage training approach, which involves generating pseudo-labels from captioning data, conducting region-text contrastive pre-training, and fine-tuning on detection data. GLIP [li2021grounded] uses a single text query for the entire image and frames detection as a phrase grounding problem. More recently, Detic [zhou2022detecting] addresses long-tail detection performance by leveraging image-level supervision. In the context of open-vocabulary detection, there may be some overlap between categories seen during training and testing. We use the term “zero-shot” when a model has never been trained on the target dataset.
@@
Fine-Tuning Foundation Models is of significant interest across many application areas [hu2021lora, zhang2023adding, gao2024clip]. Standard fine-tuning procedures employ both linear probing [chen2020simple, he2022masked, he2020momentum] and full-finetuning [wang2017growing, wu2023revisiting, kirkpatrick2017overcoming, liu2024few] to adapt models to downstream tasks. However, such methods may be suboptimal as they can be computationally expensive. Instead, recent works like CLIP-Adapter [gao2024clip] and Tip-Adapter [zhang2021tip] fine-tune CLIP using parameter-efficient methods [houlsby2019parameter, zhang2020side, jia2022visual] which optimize lightweight MLPs while keeping the encoder frozen. Similarly, inspired by the success of prefix-tuning in language models [deng2022rlprompt, jiang2020can, haviv2021bertese, gao2020making], prompt adaptation [lu2022prompt, zhu2023prompt, xing2023dual, zhou2022conditional] replaces hand-crafted prompts like "a photo of a {cls}" with learned word embeddings. CoOp [zhou2022learning] and other prompting methods [lu2022prompt, zhu2023prompt, zhou2022conditional] finetune CLIP via prefix-tuning. Different from most prior work, we investigate fine-tuning strategies for VLM-based detectors using few-shot multi-modal examples.
@@ 3.1 Foundational FSOD Benchmark
Existing FSOD benchmarks repurpose well-established datasets like PASCAL VOC [Everingham10] and COCO [Lin2014MicrosoftCC] by partitioning them into base and novel classes for pre-training and fine-tuning, respectively. For COCO, the 60 categories disjoint with PASCAL VOC are used as base classes and the remaining 20 are used as novel classes [wang2020frustratingly]. However, this setup is artificial and does not reflect how FSOD is deployed in practice. First, FSOD benchmarks construct a set of novel classes that include common concepts such as car and person, and require FSOD methods to detect these common classes by assuming they have only few examples. Importantly, VLMs like GroundingDINO [liu2023grounding] can already detect common categories with high accuracy on COCO without fine-tuning (cf. Table 1). Therefore, we focus on benchmarking Foundational FSOD on more realistic and challenging datasets like nuImages [caesar2020nuscenes]. Second, existing FSOD benchmarks require that datasets are partitioned into base and novel classes, which is infeasible for large-scale (often private) foundational datasets. For example, although CLIP’s [radford2021learning] model weights are publicly available, its pre-training dataset is not. Instead, FSOD methods should only fine-tune VLMs on $K$-shot annotations for $C$ target classes (or novel, as termed in conventional FSOD benchmarks), and also evaluate performance on these $C$ classes.
@@ 3.2 Few-Shot Multi-Modal Concept Alignment
Although VLMs achieve strong zero-shot performance on common classes, they struggle when the target class is different from concepts encountered on the web (cf. Fig. 1). For example, nuImages [caesar2020nuscenes] defines the cab of a truck as a separate concept from its trailer. However, Detic [zhou2022detecting] detects the entire vehicle as truck. This fine-grained distinction is provided to human annotators with visual examples and textual descriptions. We explore seven methods for alignment (either explicitly by updating model weights via gradient-based fine-tuning or in-context via prompting) below.
Prompt Engineering uses expressive descriptions, attributes [menon2022visual], or synonyms [parashar2023prompting, parashar2024neglected] in the text prompt to manually improve the alignment of foundation model outputs to target concepts of interest. In our case, we prompt VLMs with synonyms of the nuImages classes to improve detection accuracy. For example, we augment the language query for pushable-pullable with synonyms like cart and wheel barrow. We provide a full list of synonyms in Table 6.
@@
Multi-Modal Prompting combines language and visual prompting to leverage multi-modal features. Intuitively, multi-modal cues can yield better alignment than uni-modal cues alone; in the above case, ambiguous concepts such as debris can be clarified with both textual descriptions (e.g trash can and tree branch) and visual examples (similar to the multi-modal annotator instructions in Fig. 1!). Here, visual and language prompts are tokenized and separately fed as inputs to a frozen VLM. Specifically, MQDet [xu2024multi] introduces a lightweight Gated Class Scalable Perceiver module that fuses visual cues and text descriptions in the text encoder via class-wise cross attention layers.
@@ 4 Experiments
We conduct extensive experiments to validate that zero-shot inference from VLMs significantly improves over state-of-the-art FSOD approaches, suggesting that existing benchmarks should be re-framed to allow foundation models to “enter the conversation”. Moreover, we demonstrate that leveraging language cues, especially those available for free (e.g., class names), are crucial to improving performance on data-constrained tasks like FSOD.
Datasets and Metrics. We repurpose nuImages [caesar2020nuscenes] to support the study of Foundational FSOD. This dataset annotates 18 classes, which are divided into groups with many, medium, and few examples [peri2023towards, ma2023longtailed]. We report average precision (AP) for each cohort. Although this dataset is not traditionally used for FSOD, nuImages’ open-world categories like debris and pushable-pullable make it particularly challenging (even for VLMs), and is a realistic benchmark for Foundational FSOD. We follow the $K$-shot dataset creation process established by [wang2020frustratingly], described below. To construct a $K$-shot dataset, we select a target class $c$ and an image at random. If the total annotations for class $c$ in the image are less than or equal to $K$, we add the image to our dataset. We repeat this process for all classes until we have exactly $K$ annotations per class. Since the specific $K$ examples can have a significant impact on the overall performance, we run each experiment over three random data splits and report the average.